On the Existence of Solutions for a Class of

Nonlinear Inclusions

Shui-Hung Hou Department of Applied Mathematics The Hong Kong Polytechnic University

Abstract

Sufficient conditions for the existence of solutions for a class of nonlinear inclusions in infinite dimensional spaces are established. The results are obtained by means of fixed-point theorem.